# Bochner-type Theorems for Distributional Category

### v1

@@ 5. Extensions for $c$-symplectic manifolds
In this section, we explain how our results from Section 4 on distributional category can be refined in the setting of closed $c$-symplectic manifolds.
@@ 5.A. Manifolds with non-negative Ricci curvature
Let $M$ be a closed $c$-symplectic $2n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}=T^{2k}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
First, we note that such a splitting exists, see Theorems 3.1 and 3.3. Since $W$ is a closed simply-connected $c$-symplectic $(2n-2k)$-dimensional manifold, we have $\operatorname{{\mathsf{cup}}}(W)=\operatorname{{\mathsf{cat}}}(W)=n-k$, see Section 2.D. Thus, we get
@@
We note that without the non-negative Ricci curvature hypothesis on $c$-symplectic manifolds $M$ having $\pi_{1}(M)$ infinite, even the (relatively) weaker inequality $b_{1}(M)\leq\operatorname{{\mathsf{dcat}}}(M)$ from Theorem 4.1 fails. To see this, consider the $n$-th symmetric product $SP^{n}(M_{g})$ of the closed orientable surface $M_{g}$ of genus $g>n\geq 2$, which is $c$-symplectic but cannot support a Riemannian metric of non-negative Ricci curvature, [DDJ, Section 6]. It follows from [DDJ] that
@@
For a proof of the following classical result, we refer to [Br82, Theorem VIII.3.1].
@@
Let $\Gamma^{\prime}$ be a finite index subgroup of $\Gamma$. If $\Gamma$ is torsion-free, then $\operatorname{{\rm cd}}(\Gamma)=\operatorname{{\rm cd}}(\Gamma^{\prime})$.
@@
Let $M$ be a closed $c$-symplectic $2n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is torsion-free. If $M^{\prime}=T^{2k}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
Because of Proposition 5.1 and Theorem 2.1, we have that
$n+k=\operatorname{{\mathsf{dcat}}}(M^{\prime})=\operatorname{{\mathsf{cat}}}(M^{\prime})\leq\operatorname{{\mathsf{cat}}}(M).$
The main result of [Dr19] says that
$\operatorname{{\mathsf{cat}}}(M)\leq\frac{\dim(M)+\operatorname{{\rm cd}}(\pi_{1}(M))}{2}=\frac{2n+\operatorname{{\rm cd}}(\pi_{1}(M))}{2}.$
Since $\pi_{1}(M)$ is torsion-free and $M^{\prime}$ is a finite cover of $M$ (see Theorem 3.1), Serre’s Theorem 5.3 implies that $\operatorname{{\rm cd}}(\pi_{1}(M))=\operatorname{{\rm cd}}(\pi_{1}(M^{\prime}))$. Since $W$ is simply connected, $\pi_{1}(M^{\prime})=\pi_{1}(T^{2k})={\mathbb{Z}}^{2k}$. Hence, $\operatorname{{\rm cd}}(\pi_{1}(M))=2k$. Therefore, we get from the above two inequalities that $\operatorname{{\mathsf{cat}}}(M)=n+k$. By Proposition 5.1 and Theorem 2.2, we have
@@
Let $M$ be a closed $c$-symplectic $2n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is torsion-free. Then $\operatorname{{\rm cd}}(\pi_{1}(M))$ is even and is the rank of the torus in the Cheeger–Gromoll splitting of $M$ implied by Theorem 3.1.
@@ 5.B. The Gottlieb group
We begin by noting that from the proof of Lemma 3.5, if $M$ is a closed $c$-symplectic $2n$-manifold, $A$ is a finitely generated free abelian finite index subgroup of $\pi_{1}(M)$, and $M^{\prime}\simeq T^{r}\times Y$ is a Hurewicz splitting of $M$ corresponding to $A$, then $r=2k$ and $M^{\prime}$ is a closed $c$-symplectic $2n$-manifold.
@@
First, we note from the proof of Theorem 4.2 that $Y$ is simply connected. Furthermore, $M^{\prime}$ and $Y$ are $c$-symplectic manifolds of dimensions $2n$ and $2n-2k$, respectively, by Lemma 3.5. Thus, we have that $n-k=\operatorname{{\mathsf{cup}}}(Y)=\operatorname{{\mathsf{cat}}}(Y)$. Finally, Theorems 2.1 and 2.2 give us the following series of inequalities:
@@
If $M^{2n}=T^{2k}\times Y$ for a closed symplectic manifold $Y$ having finite $\pi_{1}(Y)$ with a trivial center, then
$\operatorname{{\mathsf{dcat}}}(M)\geq\operatorname{{\mathsf{dcat}}}(M^{\prime})=\frac{\dim(M)}{2}+k.$
@@
By [Gom95], such a $Y$ exists. We have $G(M)={\mathbb{Z}}^{k}$ since $\pi_{1}(Y)$ has trivial center, and $G(M)$ has finite index in $\pi_{1}(M)$ since $\pi_{1}(Y)$ is finite. By Proposition 5.7, we get $\operatorname{{\mathsf{dcat}}}(M^{\prime})=\operatorname{{\mathsf{cat}}}(M^{\prime})=n+k$. The standard inequality $\operatorname{{\mathsf{dcat}}}(M)\geq\operatorname{{\mathsf{dcat}}}(M^{\prime})$ completes the proof. ∎
As in Section 5.A, we can say more using Serre’s Theorem 5.3 and the main result of [Dr19] in the case when the groups are torsion-free.
@@
Because of Proposition 5.7 and Theorem 2.1, we have that
$n+k=\operatorname{{\mathsf{dcat}}}(M^{\prime})=\operatorname{{\mathsf{cat}}}(M^{\prime})\leq\operatorname{{\mathsf{cat}}}(M).$
In light of Theorem 2.2, it suffices to prove $\operatorname{{\mathsf{cat}}}(M)\leq n+k$. Proceeding along the lines of the proof of Theorem 5.4, we get
$\operatorname{{\mathsf{cat}}}(M)\leq\frac{2n+\operatorname{{\rm cd}}(\pi_{1}(M^{\prime}))}{2}$
due to Serre’s theorem and Dranishnikov’s theorem. But $M^{\prime}\simeq T^{2k}\times Y$, and $Y$ is simply connected. So, $\operatorname{{\rm cd}}(\pi_{1}(M^{\prime}))=2k$. This completes the proof. ∎
@@
in Theorem 4.1 is tantalizingly close to saying that $\operatorname{{\mathsf{cat}}}(M)=\operatorname{{\mathsf{dcat}}}(M)$ for all such $M$ (since the Cheeger–Gromoll splitting $T^{k}\times W$ implies $W\simeq\widetilde{M}$). In both these results, if $\pi_{1}(M)$ is torsion-free, then $b_{1}(M)$ may be replaced by $\operatorname{{\rm cd}}(\pi_{1}(M))$. We leave it to the reader to ponder whether $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)$ for such manifolds.
@@ 6. Distributional category and macroscopic dimension
Before we end, we would like to relate the distributional category to another very geometric type of invariant in the Bochner fashion. The notion of the macroscopic dimension of Riemannian manifolds was introduced by Gromov in [Gr96], in part to study the problem of the (non-)existence of Riemannian metrics of positive scalar curvature (see also [Dr13][DDJ]). For a Riemannian manifold $X$, its macroscopic dimension, denoted $\dim_{mc}X$, is the smallest integer $n$ for which there exists a continuous proper map $f:X\to K$ to an $n$-dimensional simplicial complex $K$ such that for some $c>0$, $\text{diam}(f^{-1}(y))<c$ for all $y\in K$.
@@
For closed manifolds $M$, a strict improvement to the above inequality was recently obtained in [DDJ, Theorem 7.18] using LS-category as follows:
@@
We further improve this inequality in the case of manifolds having non-negative Ricci curvature111 There is a different notion of macroscopic dimension due to Dranishnikov (see [Dr13]), which is defined using Lipschitz maps instead of continuous maps. For closed manifolds $M$ having non-negative Ricci curvature, these two notions coincide on $\widetilde{M}$— see [DDJ, Proposition 7.10]. using distributional category.
@@
Let $M$ be a closed $n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}=T^{r}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
The proof of [DDJ, Proposition 7.10] implies that $\dim_{mc}\widetilde{M}=r$. But we have from the proof of Theorem 4.1 that $r+\operatorname{{\mathsf{cup}}}(W)\leq\operatorname{{\mathsf{dcat}}}(M^{\prime})$. Therefore,
@@
If $M$ is flat, then $M$ is aspherical and $\dim_{mc}\widetilde{M}=\dim(M)=\operatorname{{\mathsf{dcat}}}(M)$ by [DDJ, Corollary 7.7] and Theorem 2.4, respectively. Conversely, if $\dim_{mc}\widetilde{M}=\operatorname{{\mathsf{dcat}}}(M)$, then we get from the above series of inequalities that $\operatorname{{\mathsf{cup}}}(W)=0$ for the closed simply connected manifold $W$. This means $W=\ast$ and hence, $M^{\prime}=T^{n}$ (as in the proof of Theorem 4.1). Therefore, $M$ is a closed aspherical $n$-manifold and $\pi_{1}(M)$ is torsion-free. Since $\pi_{1}(M^{\prime})={\mathbb{Z}}^{n}$ has finite index in $\pi_{1}(M)$, the generalization of Vasquez’s characterization [Va70] mentioned in Section 4 implies that there exists a closed flat $n$-manifold $N=K(\pi_{1}(M),1)$. Since $M=K(\pi_{1}(M),1)$, we must have $M\simeq N$. In the cases $n\neq 3,4$, the results of Farrell–Hsiang [FH83] imply that the corresponding surgery structure set is zero and so $M=N$ (see Section 4). For $n=3,4$[KL09, Theorems 0.7 & 0.11] imply that the flat manifold $N$ is strongly Borel, which means the homotopy equivalence $N\to M$ above is homotopic to a homeomorphism. Therefore, in any case, $M$ is flat. ∎
@@
The fact that $\dim_{mc}\widetilde{M}=\dim(M)$ implies the flatness of $M$ was proved in [DDJ, Proposition 7.10] in a completely different way, so once we showed in the proof above that $M$ is aspherical, we could have simply applied that result along with Theorem 2.4. Nevertheless, it seems worthwhile to see how a more classical approach leads to the flatness conclusion.
@@
Observe that $\dim_{mc}\widetilde{M}\leq\dim_{mc}\widetilde{N}+\dim_{mc}\widetilde{Y}=\dim(N)$ because $N$ is aspherical and $\widetilde{Y}$ is compact. Then by definition of macroscopic dimension, we must have $\dim_{mc}\widetilde{M}=\dim(N)$. Now, since $M$ is not flat, Theorem 6.2 implies $\dim(N)=\dim_{mc}\widetilde{M}<\operatorname{{\mathsf{dcat}}}(M)$. ∎
@@
The hypothesis of non-negative Ricci curvature in Theorem 6.2 is necessary, even for the weaker implication that $\dim_{mc}\widetilde{M}=\operatorname{{\mathsf{dcat}}}(M)\implies M$ is aspherical. To see this, consider the $g$-th symmetric product $SP^{g}(M_{g})$ of the closed orientable surface $M_{g}$ of genus $g\geq 2$ and note its following properties from [DDJ]: $SP^{g}(M_{g})$ cannot support a metric of non-negative Ricci curvature, $\dim_{mc}\widetilde{SP^{g}(M_{g})}=2g=\operatorname{{\mathsf{dcat}}}(SP^{g}(M_{g}))$, and $\pi_{2}(SP^{g}(M_{g}))$ is infinite.
@@
We conclude by pointing out that the arguments of Proposition 5.1 and Theorem 6.2 give us the improved bound $\dim_{mc}\widetilde{M}\leq 2\operatorname{{\mathsf{dcat}}}(M)-\dim(M)$ for any closed $c$-symplectic manifold $M$ having non-negative Ricci curvature and infinite fundamental group. However, the equality $\dim_{mc}\widetilde{M}=2\operatorname{{\mathsf{dcat}}}(M)-\dim(M)$ here does not imply that $M$ is flat (or even aspherical). To see this, consider $M=T^{2k}\times N$ for a closed simply connected symplectic $(2n-2k)$ manifold $N$ with non-negative Ricci curvature. Then, $\dim_{mc}\widetilde{M}=2k$ and $\operatorname{{\mathsf{dcat}}}(M)=n+k$, but $M$ is not aspherical. The same example also shows that in the setting of Proposition 5.1, the equality $b_{1}(M)=2\operatorname{{\mathsf{dcat}}}(M)-\dim(M)$ does not imply that $M$ is a torus.

### v2

@@ 5. Extensions for $c$-symplectic manifolds
Our results from Section 4 illustrate that the techniques used to lower bound LS-category (see [Opr02a][Opr02b]) by $b_{1}(M)$ and $\operatorname{{\mathsf{rank}}}(G(X))$ remain useful in lower bounding distributional category as well. So, it is natural to ask whether $\operatorname{{\mathsf{dcat}}}=\operatorname{{\mathsf{cat}}}$ in the settings of Theorems 4.1 and 4.2. In this section, we address this interesting question by explaining how our results from Section 4 can be refined in the setting of closed $c$-symplectic manifolds and $\operatorname{{\mathsf{dcat}}}=\operatorname{{\mathsf{cat}}}$ can be obtained.
@@ 5.A. Manifolds with non-negative Ricci curvature
Let $M$ be a closed $c$-symplectic Riemannian $2n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}\cong T^{2k}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
First, we note that such a splitting exists, see Theorems 3.1 and $3.$ Since $W$ is a closed simply-connected $c$-symplectic (2n2k)(2n-2k)-dimensional manifold, we have $\operatorname{{\mathsf{cup}}}(W)=\operatorname{{\mathsf{cat}}}(W)=n-k$, see Section 2.D. Thus, we get
@@
We note that without the non-negative Ricci curvature hypothesis on $c$-symplectic manifolds $M$ having infinite fundamental group, even the (relatively) weaker inequality $b_{1}(M)\leq\operatorname{{\mathsf{dcat}}}(M)$ from Theorem 4.1 fails. To see this, consider the $n$-th symmetric product $SP^{n}(M_{g})$ of the closed orientable surface $M_{g}$ of genus $g>n\geq 2$, which is $c$-symplectic but cannot support a Riemannian metric of non-negative Ricci curvature, [DCDJ, Section 5]. It follows from [DCDJ, Section 6] that
@@
For a proof of the following classical result, we refer to [Bro82, Theorem VIII.3.1].
@@
Let $\Gamma^{\prime}$ be a finite index subgroup of $\Gamma$. If $\Gamma$ is torsion-free, then $\operatorname{{\mathsf{cd}}}(\Gamma)=\operatorname{{\mathsf{cd}}}(\Gamma^{\prime})$.
@@
Let $M$ be a closed $c$-symplectic Riemannian $2n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is torsion-free. If $M^{\prime}\cong T^{2k}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
Because of Proposition 5.1 and Theorem 2.1 above, we have the inequality $n+k=\operatorname{{\mathsf{dcat}}}(M^{\prime})=\operatorname{{\mathsf{cat}}}(M^{\prime})\leq\operatorname{{\mathsf{cat}}}(M)$. The main result of [Dra19] says that
$\operatorname{{\mathsf{cat}}}(M)\leq\frac{\dim(M)+\operatorname{{\mathsf{cd}}}(\pi_{1}(M))}{2}=\frac{2n+\operatorname{{\mathsf{cd}}}(\pi_{1}(M))}{2}.$
Since $\pi_{1}(M)$ is torsion-free and $M^{\prime}$ is a finite cover of $M$ (see Theorem 3.1), Serre’s Theorem 5.3 implies that $\operatorname{{\mathsf{cd}}}(\pi_{1}(M))=\operatorname{{\mathsf{cd}}}(\pi_{1}(M^{\prime}))$. Since $W$ is simply connected, $\pi_{1}(M^{\prime})\cong\pi_{1}(T^{2k})\cong{\mathbb{Z}}^{2k}$. Hence, $\operatorname{{\mathsf{cd}}}(\pi_{1}(M))=2k$. Therefore, we get from the above two inequalities that $\operatorname{{\mathsf{cat}}}(M)=n+k$. By Proposition 5.1 and Theorem 2.2, we have
@@
Let $M$ be a closed $c$-symplectic $2n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is torsion-free. Then $\operatorname{{\mathsf{cd}}}(\pi_{1}(M))$ is even and is the rank of the torus in the Cheeger–Gromoll splitting of $M$ implied by Theorem 3.1.
@@ 5.B. The Gottlieb group
We begin by noting that from the proof of Lemma 3.5, if $M$ is a closed $c$-symplectic $2n$-manifold, AA is a finitely generated free abelian finite index subgroup of $\pi_{1}(M)$, and $M^{\prime}\simeq T^{r}\times Y$ is a Hurewicz splitting of $M$ corresponding to $A$, then $r=2k$ and $M^{\prime}$ is a closed $c$-symplectic $2n$-manifold.
@@
First, we note from the proof of Theorem 4.2 that $Y$ is simply connected. Furthermore, $M^{\prime}$ and $Y$ are $c$-symplectic manifolds of dimensions $2n$ and 2n2k2n-2k, respectively, by Lemma 3.5. Thus, we have that $n-k=\operatorname{{\mathsf{cup}}}(Y)=\operatorname{{\mathsf{cat}}}(Y)$. Finally, Theorems 2.1 and 2.2 give us the following series of inequalities:
@@
If $M^{2n}=T^{2k}\times Y$ for a closed symplectic manifold $Y$ having finite $\pi_{1}(Y)$ with a trivial center, then $\operatorname{{\mathsf{dcat}}}(M)\geq n+k$.
@@
By [Gom95], such a manifold $Y$ exists. We have $G(M)\cong{\mathbb{Z}}^{k}$ since $\pi_{1}(Y)$ has trivial center, and $G(M)$ has finite index in $\pi_{1}(M)$ since $\pi_{1}(Y)$ is finite. By Proposition 5.7, we get $\operatorname{{\mathsf{dcat}}}(M^{\prime})=\operatorname{{\mathsf{cat}}}(M^{\prime})=n+k$. The standard inequality $\operatorname{{\mathsf{dcat}}}(M)\geq\operatorname{{\mathsf{dcat}}}(M^{\prime})$ completes the proof. ∎
As in Section 5.A, we can say more using Serre’s Theorem 5.3 and the main result of [Dra19] in the case when the groups are torsion-free.
@@
Because of Proposition 5.7 and Theorem 2.1 above, we have the inequality $n+k=\operatorname{{\mathsf{dcat}}}(M^{\prime})=\operatorname{{\mathsf{cat}}}(M^{\prime})\leq\operatorname{{\mathsf{cat}}}(M)$. In light of Theorem 2.2, it suffices to prove $\operatorname{{\mathsf{cat}}}(M)\leq n+k$. Proceeding along the lines of the proof of Theorem 5.4, we get
$\operatorname{{\mathsf{cat}}}(M)\leq\frac{2n+\operatorname{{\mathsf{cd}}}(\pi_{1}(M^{\prime}))}{2}$
due to Serre’s theorem and Dranishnikov’s theorem. But $M^{\prime}\simeq T^{2k}\times Y$, and $Y$ is simply connected. So, $\operatorname{{\mathsf{cd}}}(\pi_{1}(M^{\prime}))=2k$. This completes the proof. ∎
@@
in Theorem 4.1 is tantalizingly close to saying that $\operatorname{{\mathsf{cat}}}(M)=\operatorname{{\mathsf{dcat}}}(M)$ for all such $M$ (since the Cheeger–Gromoll splitting $T^{k}\times W$ implies $W\simeq\widetilde{M}$). In both these results, if $\pi_{1}(M)$ is torsion-free, then $b_{1}(M)$ may be replaced by $\operatorname{{\mathsf{cd}}}(\pi_{1}(M))$. We leave it to the reader to ponder whether $\operatorname{{\mathsf{dcat}}}(M)=\operatorname{{\mathsf{cat}}}(M)$ for such manifolds.
@@ 6. Distributional category and macroscopic dimension
Before we end, we would like to relate the distributional category to another very geometric type of invariant in the Bochner fashion. The notion of the macroscopic dimension of Riemannian manifolds was introduced by Gromov in [Gro96], in part to study the problem of the (non-)existence of Riemannian metrics of positive scalar curvature (see also [Dra13][DCDJ]). For a Riemannian manifold $X$, its macroscopic dimension, denoted $\dim_{mc}X$, is the smallest integer $n$ for which there exists a continuous proper map $f:X\to K$ to an $n$-dimensional simplicial complex $K$ such that for some $c>0$, $\text{diam}(f^{-1}(y))<c$ for all $y\in K$.
@@
For closed manifolds $M$, a strict improvement to the above inequality was recently obtained in [DCDJ, Theorem 7.18] using LS-category as follows:
@@
We further improve this inequality in the case of manifolds having non-negative Ricci curvature222 There is a different notion of macroscopic dimension due to Dranishnikov (see [Dra13]), which is defined using Lipschitz maps instead of continuous maps. For closed manifolds $M$ having non-negative Ricci curvature, these two notions coincide on $\widetilde{M}$— see [DCDJ, Proposition 7.10]. using distributional category.
@@
Let $M$ be a closed Riemannian $n$-manifold with non-negative Ricci curvature such that $\pi_{1}(M)$ is infinite. If $M^{\prime}\cong T^{r}\times W$ is a Cheeger–Gromoll splitting of $M$, then
@@
The proof of [DCDJ, Proposition 7.10] implies that $\dim_{mc}\widetilde{M}=r$. But we have from the proof of Theorem 4.1 that $r+\operatorname{{\mathsf{cup}}}(W)\leq\operatorname{{\mathsf{dcat}}}(M^{\prime})$. Therefore,
@@
If $M$ is flat, then $M$ is aspherical and $\dim_{mc}\widetilde{M}=\dim(M)=\operatorname{{\mathsf{dcat}}}(M)$ by [DCDJ, Corollary 7.7] and Theorem 2.4, respectively. Conversely, if $\dim_{mc}\widetilde{M}=\operatorname{{\mathsf{dcat}}}(M)$, then we get from the above series of inequalities that $\operatorname{{\mathsf{cup}}}(W)=0$ for the closed simply connected manifold $W$. This means $W\simeq\ast$ and hence, $M^{\prime}\cong T^{n}$ (as in the proof of Theorem 4.1). Therefore, $M$ is a closed aspherical $n$-manifold and $\pi_{1}(M)$ is torsion-free. Since $\pi_{1}(M^{\prime})\cong{\mathbb{Z}}^{n}$ has finite index in $\pi_{1}(M)$, the generalization of Vasquez’s characterization [Vas70] mentioned in Section 4 implies that there exists a closed flat $n$-manifold $N=K(\pi_{1}(M),1)$. Since $M=K(\pi_{1}(M),1)$, we must have $M\simeq N$. In the cases $n\neq 3,4$, the results of Farrell–Hsiang [FH83] imply that the corresponding surgery structure set is zero and so $M\cong N$ (see Section 4). For n=3,4n=3,4[KL09, Theorems 0.7 & 0.11] imply that the flat manifold $N$ is strongly Borel, which means the homotopy equivalence $N\to M$ above is homotopic to a homeomorphism. Therefore, in any case, $M$ is flat. ∎
@@
The fact that $\dim_{mc}\widetilde{M}=\dim(M)$ implies the flatness of $M$ was proved in [DCDJ, Proposition 7.10] in a completely different way, so once we showed in the proof above that $M$ is aspherical, we could have simply applied that result along with Theorem 2.4. Nevertheless, it seems worthwhile to see how a more classical approach leads to the flatness conclusion.
@@
Observe that $\dim_{mc}\widetilde{M}\leq\dim_{mc}\widetilde{N}+\dim_{mc}\widetilde{Y}=\dim(N)$ because $N$ is aspherical and $\widetilde{Y}$ is compact. Then by the definition of macroscopic dimension, we must have $\dim_{mc}\widetilde{M}=\dim(N)$. But because $M$ is not flat, Theorem 6.2 implies $\dim(N)=\dim_{mc}\widetilde{M}<\operatorname{{\mathsf{dcat}}}(M)$. ∎
@@
The hypothesis of non-negative Ricci curvature in Theorem 6.2 is necessary. To see this, consider the $g$-th symmetric product $SP^{g}(M_{g})$ of the closed orientable surface $M_{g}$ of genus $g\geq 2$ and note its following properties from [DCDJ]: $SP^{g}(M_{g})$ cannot support a metric of non-negative Ricci curvature, $\dim_{mc}\widetilde{SP^{g}(M_{g})}=2g=\operatorname{{\mathsf{dcat}}}(SP^{g}(M_{g}))$, and $\pi_{2}(SP^{g}(M_{g}))$ is infinite.
@@
We conclude by pointing out that the arguments of Proposition 5.1 and Theorem 6.2 give us the improved bound $\dim_{mc}\widetilde{M}\leq 2\operatorname{{\mathsf{dcat}}}(M)-\dim(M)$ for any closed $c$-symplectic manifold $M$ having non-negative Ricci curvature and infinite fundamental group. However, the equality $\dim_{mc}\widetilde{M}=2\operatorname{{\mathsf{dcat}}}(M)-\dim(M)$ here does not imply that $M$ is flat (or even aspherical). To see this, consider $M=T^{2k}\times N$ for a closed simply connected symplectic (2n2k)(2n-2k) manifold $N$ with non-negative Ricci curvature. Then, $\dim_{mc}\widetilde{M}=2k$ and $\operatorname{{\mathsf{dcat}}}(M)=n+k$, but $M$ is not aspherical. The same example also shows that in the setting of Proposition 5.1, the equality $b_{1}(M)=2\operatorname{{\mathsf{dcat}}}(M)-\dim(M)$ does not imply that $M$ is a torus.
Acknowledgement
The authors thank the anonymous referee for valuable suggestions.